A pen shop sells pen for 8php each and sold 250 pens per day on average. They estimate that the price elasticity of demand for pen is (-) 1.6. Calculate the expected number of customers if they reduce the pen price to 7php each.



1
Expert's answer
2021-09-14T09:31:36-0400

Price elasticity of demand = "\frac{Percentage change in quantity demanded}{Percentage change in price}"


Change in price "= 8-7 = 1"


Percentage change in price = "\frac {1}{8}× 100 =" 12.5%


When price elasticity of demand = (-) 1.6, the pen sales increases by "1.6 \u00d7 12.5 = 20" %


= "\frac{20}{100} ×250 = 50" extra pens sold


Expected number of customers "=250+50 = 300"